# Spin conductances and magnetization production in chiral molecular junctions

### v2

@@ Abstract
Motivated by experimental reports on chirality induced spin selectivity, we investigate a minimal model that allows us to calculate the charge and spin conductances through helical molecules analytically. The spin-orbit interaction is assumed to be non-vanishing on the molecule and negligible in the reservoirs (leads). The band-structure of the molecule features four helical modes with spin-momentum locking that are analogous of edge-currents in the quantum spin Hall effect. While charge is conserved and therefore the charge current is independent of where it is measured, - reservoirs or molecule, - our detailed calculations reveal that the spin currents in the left and right lead are equal in magnitudes but with opposite signs (in linear response). We predict that transport currents flowing through helical molecules are accompanied by a spin accumulation in the contact region with the same magnetization direction for source and drain. Further, we predict that the spin-conductance can be extracted directly from measuring the (quasi-static) spin accumulation - rather than the spin current itself, which is very challenging to obtain experimentally.
keywords:
@@ 1 Introduction
Charge currents are routinely measured and analyzed in molecular electronics 1. The discovery of a family of phenomena that exhibit chirality induced spin selectivity (CISS) has led to a resurge of interest in spin-related phenomena in this field.2, 3, 4, 5, 6, 7 Specifically, experiments report a strong correlation between molecular chirality and a preferred spin direction in systems that exhibit a (nominally) very weak spin-orbit interaction. At present, there is no consensus concerning the explanation of many experimental CISS results 8.
@@
Motivated by the CISS phenomena, we here address charge and spin currents in chiral molecular junctions within the framework of a minimal model. As a diagnostic tool of the junction’s atomic structure, spin currents offer advantages as compared to charge currents: Spin polarized currents can, in principle, be detected in charge-transport measurements within an analyzer-polarizer setup employing, e.g., magnetized leads. However, they are expected to manifest themselves only in the non-linear regime and not in the linear charge conductance $G(\mathbf{M})$. This is a consequence of Onsager’s reciprocity; its importance for the theory of the CISS effect was emphasized by Yang, van der Wal and van Wees9, 10. iiiAs is well known, due to the Onsager relations magnetotransport experiments based on the polarizer-analyzer setup will not detect spin-currents in the linear transport regime: even if a non-vanishing spin-current is flowing, the charge conductance is independent of the magnetization direction $G(\mathbf{M})=G(-\mathbf{M})$ and therefore insensitive to the flow of spin. The spin-conductance, on the other hand, is less restricted by Onsager symmetries. It can be inferred, at least in principle, from measuring the pile-up of magnetization, e.g., in source or drain. A brief review of the symmetry properties of transport coefficients is available in the Supporting Information, Sec. S1.
@@
These considerations motivated us and other researchers11 to investigate a minimal model for a chiral molecule in the presence of spin-orbit interactions. While Ref.11 has focussed on the effect of contact-enhanced SOI, we adopt the model proposed by Michaeli and Naaman12; it exhibits SOI on the entire molecule and has the extra benefit of allowing for an analytical treatment. While Michaeli and Naaman have studied the transmission properties, our focus is on (spin) conductances. As one would have expected, the spin conductance turns out to be non-vanishing due to spin-orbit coupling. As a consequence, the transmitted and reflected currents tend to build up a non-vanishing spin-accumulation near both contacts, source and drain, already in the linear regime13. In stationary non-equilibrium the magnitude of induced magnetization is likely controlled by spin-relaxation processes. We predict the orientation of the magnetization in both contacts to be the same, in agreement with requirements of time-reversal invariance (TRI).11
@@
Our work is of potential impact for constructing a molecular machine. Since total angular momentum is conserved by the $LS$-coupling, spin-flip processes exert a mechanical torque that can drive an engine. An analogue driving mechanism based on angular transfer has been investigated in Ref. 14.
2 Results
2.1 Minimal modeling of a helical molecule
@@
Following Michaeli and Naaman12 we consider electrons bound to a long tube. The left and right side of the tube are of cylindrical shape and represent (semi-infinite) reservoirs; the central region in between takes a helical shape to mimic a chiral molecule, see Fig. 1.
@@
In the limit of small $d/\tilde{R}$ the electronic wavefunction is tightly bound to the helix and the quasi one-dimensional nature of the model becomes manifest; here, $\tilde{R}=\sqrt{(2\pi R)^{2}+b^{2}}>0$ denotes the distance covered when completing one helical turn. In this limit, the longitudinal and transverse motion approximately decouple and the problem simplifies. Following this idea, a systematic expansion of Eq. (1) in $d/\tilde{R}$ has been performed by Michaeli and Naaman12 and keeping only leading order terms a minimal model has been derived.
Refer to captionRefer to caption
Figure 1: Top: Scheme of the molecular junction consisting of a helical tube (center) and a pair of semi-infinite straight tubes, left and right. The radius $R$, pitch $b$, thickness $2d$ are indicated. Bottom: Illustration of helical coordinates $s,\varrho,\theta$: Green line shows part of a helix that evolves along the $z$ axis (dash-dotted line) with unit vector $\hat{\mathbf{z}}$. At any given point on the helix, the unit vector $\hat{\mathbf{n}}$ is perpendicular to $\hat{\mathbf{z}}$ and the unit tangent vector $\hat{\mathbf{s}}$. In helical coordinates, $s$ is the distance along the helix; $\theta$ and $\varrho$ are polar coordinates in a plane normal to $\hat{\mathbf{s}}$; $\theta$ measures from $\hat{\mathbf{n}}$.
2.2 Limit of a narrow tube
2.2.1 Wavefunction factorization
@@
For simplicity, we will assume local rotational invariance in the sense $V_{\mathrm{H}}(\mathbf{r})=V_{\mathrm{H}}(\varrho)$. In the small $d/\tilde{R}$ limit, the longitudinal and transverse motion nearly decouple,12 and the wavefunction factorizes,iiiiiiThe details of the confinement $V_{\mathrm{H}}(\varrho)$ are not relevant for the main conclusions of our work. Focusing on the lowest-energy eigenstates, we adopt a harmonic approximation $V_{\mathrm{H}}(\varrho)\approx\frac{1}{2}\frac{\hbar^{2}}{m_{\mathrm{e}}d^{4}}\varrho^{2}$: the resulting wavefunctions $\Phi_{N,m}(\theta,\varrho)$ are eigenstates of a 2D harmonic oscillator with $m$-independent eigen-energies $E_{N}$ and $m$ restricted to $-N,-N+2,\ldots,N-2,N$. $\Psi(s,\varrho,\theta)\approx\bar{\Psi}_{m}(s)\Phi_{N,m}(\varrho)\,e^{\mathrm{i}m\,\theta}$. The quantum number $N\in\mathbb{N}_{0}$ governs the nodal structure of the wavefunction in the radial direction. Due to the stipulated rotational invariance, the $\hat{\mathbf{s}}$-component of the angular momentum, $-\mathrm{i}\hbar\partial/\partial\theta$, is a good quantum number; we call it $\hbar m$ and $m\in\mathbb{Z}$. Finally, owing to the presence of SOI, $\bar{\Psi}_{m}(s)$ represents a two-spinor in spin-space.
@@
2.2.2 Spin-orbit coupling term
In the narrow-tube limit an expression for the spin-orbit term was given in the Ref. 12,
@@ 2.2.2 Spin-orbit coupling term
As $s$ increases along the helix, $\mathbf{L}(s)$ precesses around the $z$-direction. This spatial dependency of ${\mathbf{L}}(s)$ is the main entry point of helicity into the quasi-one-dimensional model. The precession of ${\mathbf{L}}$ invites an analogy with a magnetic moment ${\bf\it\mu}$ in a rotating magnetic field with “Zeeman energy” $|{\bf\it\mu}\cdot\mathbf{B}|=\hbar\beta m$ and ${\bf B}=B\quantity[\sin(\Omega s)\hat{\mathbf{x}}-\cos(\Omega s)\hat{\mathbf{y}}+B_{z}/B\,\hat{\mathbf{z}}]$; $s$ being the effective “time” and oscillation frequency $\Omega=2\pi/\lambda\tilde{R}$. Motivated by this observation, in the following paragraph we adopt a transformation to the rotating frame 15.
2.2.3 Non-Abelian gauge transformation
@@
2.2.4 Minimal model Hamiltonian
@@
The $\tilde{\gamma}$-term corresponds to a momentum shift and thus represents a ”synthetic” vector potential. It can formally be removed from $\hat{H}^{\prime}_{N,m}$ by dressing the wavefunctions with a gauge-factor $e^{\mathrm{i}\tilde{\gamma}ms^{\prime}}$, 16 and therefore leaves the spectrum invariant. If $\tilde{\gamma}m$ has a natural interpretation as a vector potential, the term proportional to $b/R$ in the second line of (9) is the corresponding ”synthetic Zeeman term”. Notice that vector potential, $\tilde{\gamma}m$, and the Zeeman term change sign under $m\mapsto-m$; they do not break TRI because the full Hamiltonian sums over all $m$.
2.3 Dispersion relation
@@
The model (10) exhibits a two-band dispersion $E_{m,\alpha}(k)$, where $\alpha=\pm 1$; see Supporting Information, Sec. S2, for explicit expressions.
@@
2.4 Spin and charge transport
2.4.1 Molecule bound to straight tubes
@@ 2.4.2 Basic definitions
2.4.3 Transmissions and Landauer Formulæ
The Landauer formalism relates conductances to the spectral transmission probabilities: for the transmission from $\mathfrak{R}$ to $\mathfrak{L}$, $T^{(\mathfrak{L}\mathfrak{R})}_{\sigma\sigma^{\prime},m}(E)$, and vice versa, $T^{(\mathfrak{R}\mathfrak{L})}_{\sigma\sigma^{\prime},m}(E)$, and to the corresponding reflection amplitudes, $R^{(\mathfrak{L}\mathfrak{L})}_{\sigma\sigma^{\prime},m}(E)$ and $R^{(\mathfrak{R}\mathfrak{R})}_{\sigma\sigma^{\prime},m}(E)$; see Ref. 21 for an overview. Here, the right (left) subscripts and superscripts of the transmission probability label the quantum numbers of an incoming (outgoing) wave, respectively. For example, $T^{(\mathfrak{R}\mathfrak{L})}_{\uparrow\downarrow,m}$ denotes the probability to transmit an electron with spin down from the left lead to the right one while flipping its spin in eigenchannel $m$. The charge conductances and the $z$-components of the spin conductances are given by
@@ 2.4.3 Transmissions and Landauer Formulæ
where $E_{\mathrm{F}}$ indicates the Fermi energy (the Supporting Information, Sec. S5, offers a standard derivation within the scattering formalism). The expression for $G$ chosen here emphasizes transmission of all spin species from left to right. From that expression the right spin conductance is obtained by multiplication by $\frac{1}{e}\frac{\hbar}{2}\sigma$ in the right lead. The left spin conductance expressed in the Eq. (12c) can then be understood as due to the reflected flux in the $\mathfrak{L}$-lead. Particle conservation (unitarity, see Supporting Information, Sec. S4), leads to the equivalent form, Eq. (12d).
2.4.4 Transport Results
We address the transport problem by calculating the scattering matrix using conventional wavefunction matching. In this process, angular momentum $m$ matches at the two interfaces: it is conserved in the scattering process; for further details see Supporting Information.
@@ 2.4.4 Transport Results
  • Importantly, the spin conductances are non-zero for energies even far from the spin-orbit gap, where the oscillations peak at $\approx 0.2\,\frac{e}{4\pi}$. It can be shown that that $G_{x}^{(\mathfrak{L},\mathfrak{R})}(E_{\mathrm{F}})=0$ for any $E_{\mathrm{F}}$ since the Bloch functions have zero average $\hat{\sigma}_{x}$. Moreover, $G_{y}^{(\mathfrak{L},\mathfrak{R})}=0$ because the expectation values of $\hat{\sigma}_{y}$ change sign along with the sign change of $m$ (see Fig. 2), while the lead Hamiltonian is $m$-independent. In other words, the $y$ components of outgoing waves exactly cancel upon the summation over $m$.
  • Remarkably, the $\mathfrak{L}$ and $\mathfrak{R}$ spin conductances differ only up to a sign. By combining TRI and left-right reflection it is possible to derive relations between scattering matrix elements that lead to the exact identity $G_{z}^{(\mathfrak{R})}=-G_{z}^{(\mathfrak{L})}$ at any energy. (See Supporting Information, Sec. S4,S5 for the details of this symmetry analysis.) In realistic molecular junctions a symmetry of couplings to the left and right leads can not be expected. We show in the next section that in this more general situation, the magnitude of $G_{z}^{(\mathfrak{R})}$ and $G_{z}^{(\mathfrak{L})}$ no longer is the same, while the sign is still opposite due to TRI.
Figure 3: Charge conductance and the $z$-components of the spin conductances in the Michaeli model as a function of the Fermi energy $E_{\mathrm{F}}$. Solid (dashed) lines are for a helix with 3 (dashed) and 6 (solid) turns. Parameters are the same as in Fig. 2, with a spin-orbit gap between 0.4 and 0.6.
2.5 Qualitative understanding of spin transport
2.5.1 A frequent misconception

### v4

@@ Abstract
Motivated by experimental reports on chirality induced spin selectivity, we investigate a minimal model that allows us to calculate the charge and spin conductances through helical molecules analytically. The spin-orbit interaction is assumed to be non-vanishing on the molecule and negligible in the reservoirs (leads). The band-structure of the molecule features four helical modes with spin-momentum locking that are analogous of edge-currents in the quantum spin Hall effect. While charge is conserved and therefore the charge current is independent of where it is measured, reservoirs or molecule, our detailed calculations reveal that the spin currents in the left and right leads are equal in magnitudes but with opposite signs (in linear response). We predict that transport currents flowing through helical molecules are accompanied by a spin accumulation in the contact region with the same magnetization direction for source and drain. Furthermore, we predict that the spin-conductance can be extracted directly from measuring the (quasi-static) spin accumulation - rather than the spin current itself, which is very challenging to obtain experimentally.
@@ I Introduction
Charge currents are routinely measured and analyzed in molecular electronics Evers et al. (2020). The discovery of a family of phenomena that exhibit chirality induced spin selectivity (CISS) has led to a resurge of interest in spin-related phenomena in this field.Göhler et al. (2011); Naaman, Paltiel, and Waldeck (2019); Liu et al. (2020); Safari et al. (2023a, b); Theiler et al. (2023) Specifically, experiments report a strong correlation between molecular chirality and a preferred spin direction in systems that exhibit a (nominally) very weak spin-orbit interaction. At present, there is no consensus concerning the explanation of many experimental CISS results Evers et al. (2022).
@@
Motivated by the CISS phenomena, we here address charge and spin currents in chiral molecular junctions within the framework of a minimal model. As a diagnostic tool of the junction’s atomic structure, spin currents offer advantages as compared to charge currents: Spin polarized currents can, in principle, be detected in charge-transport measurements within an analyzer-polarizer setup employing, e.g., magnetized leads. However, they are expected to manifest themselves only in the non-linear regime and not in the linear charge conductance $G(\mathbf{M})$. This is a consequence of Onsager’s reciprocity; its importance for the theory of the CISS effect was emphasized by Yang, van der Wal and van WeesYang, van der Wal, and van Wees (2019, 2020). iiiAs is well known, due to the Onsager relations magnetotransport experiments based on the polarizer-analyzer setup will not detect spin-currents in the linear transport regime: even if a non-vanishing spin-current is flowing, the charge conductance is independent of the magnetization direction $G(\mathbf{M})=G(-\mathbf{M})$ and therefore insensitive to the flow of spin. The spin-conductance, on the other hand, is less restricted by Onsager symmetries. It can be inferred, at least in principle, from measuring the pile-up of magnetization, e.g., in source or drain. A brief review of the symmetry properties of transport coefficients is available in the supplementary material, Sec. S1.
@@
These considerations motivated us and other researchersWolf et al. (2022) to investigate a minimal model for a chiral molecule in the presence of spin-orbit interactions. While Ref.Wolf et al. (2022) has focussed on the effect of contact-enhanced SOI, we adopt the model proposed by Michaeli and NaamanMichaeli and Naaman (2019); it exhibits SOI on the entire molecule and has the extra benefit of allowing for an analytical treatment. While Michaeli and Naaman have studied the transmission properties, our focus is on (spin) conductances. As one would have expected, the spin conductance turns out to be non-vanishing due to spin-orbit coupling. As a consequence, the transmitted and reflected currents tend to build up a non-vanishing spin-accumulation near both contacts, source and drain, already in the linear regimeGarcía-Blázquez, Dednam, and Palacios (2023). In stationary non-equilibrium the magnitude of induced magnetization is likely controlled by spin-relaxation processes. We predict the orientation of the magnetization in both contacts to be the same, in agreement with requirements of time-reversal invariance (TRI).Wolf et al. (2022)
@@
Our work is of potential impact for constructing a molecular machine. Since total angular momentum is conserved by the $LS$-coupling, spin-flip processes exert a mechanical torque that can drive an engine. An analogue driving mechanism based on angular transfer has been investigated in Ref. Korytár and Evers (2023).
II Results
II.1 Minimal modeling of a helical molecule
@@
Following Michaeli and NaamanMichaeli and Naaman (2019) we consider electrons bound to a long tube. The left and right side of the tube are of cylindrical shape and represent (semi-infinite) reservoirs; the central region in between takes a helical shape to mimic a chiral molecule, see Fig. 1.
@@
In the limit of small $d/\tilde{R}$ the electronic wavefunction in the central region is tightly bound to the helix and the quasi one-dimensional nature of the model becomes manifest; here, $\tilde{R}=\sqrt{(2\pi R)^{2}+b^{2}}>0$ denotes the distance covered when completing one helical turn. In this limit, the longitudinal and transverse motion approximately decouple and the problem simplifies. Following this idea, a systematic expansion of Eq. (1) in $d/\tilde{R}$ has been performed by Michaeli and NaamanMichaeli and Naaman (2019) and keeping only leading order terms a minimal model has been derived.
Figure 1: Top: Scheme of the molecular junction consisting of a helical tube (center) and a pair of semi-infinite straight tubes, left and right. The radius $R$, pitch $b$, thickness $2d$ are indicated. Bottom: Illustration of helical coordinates $s,\varrho,\theta$: Green line shows part of a helix that evolves along the $z$ axis (dash-dotted line) with unit vector $\hat{\mathbf{z}}$. At any given point on the helix, the unit vector $\hat{\mathbf{n}}$ , the unit tangent vector $\hat{\mathbf{s}}$ and $\hat{\mathbf{z}}$, are perpendicular (indicated by the small rectangle) . In helical coordinates, $s$ is the distance along the helix; $\theta$ and $\varrho$ are polar coordinates of a vector $\mathbf{r}$ lying in a plane normal to $\hat{\mathbf{s}}$; $\theta$ measures from $\hat{\mathbf{n}}$.
II.2 Limit of a narrow tube
II.2.1 Wavefunction factorization
@@
For simplicity, we will assume local rotational invariance in the sense $V_{\mathrm{H}}(\mathbf{r})=V_{\mathrm{H}}(\varrho)$. In the small $d/\tilde{R}$ limit, the longitudinal and transverse motion nearly decouple,Michaeli and Naaman (2019) and the wavefunction factorizes,iiiiiiThe details of the confinement $V_{\mathrm{H}}(\varrho)$ are not relevant for the main conclusions of our work. Focusing on the lowest-energy eigenstates, we adopt a harmonic approximation $V_{\mathrm{H}}(\varrho)\approx\frac{1}{2}\frac{\hbar^{2}}{m_{\mathrm{e}}d^{4}}\varrho^{2}$: the resulting wavefunctions $\Phi_{N,m}(\theta,\varrho)$ are eigenstates of a 2D harmonic oscillator with $m$-independent eigen-energies $E_{N}$ and $m$ restricted to $-N,-N+2,\ldots,N-2,N$. $\Psi(s,\varrho,\theta)\approx\bar{\Psi}_{m}(s)\Phi_{N,m}(\varrho)\,e^{\mathrm{i}m\,\theta}$. The quantum number $N\in\mathbb{N}_{0}$ governs the nodal structure of the wavefunction in the radial direction. Due to the stipulated rotational invariance, the $\hat{\mathbf{s}}$-component of the angular momentum, $-\mathrm{i}\hbar\partial/\partial\theta$, is a good quantum number; we call it $\hbar m$ and $m\in\mathbb{Z}$. Finally, owing to the presence of SOI, $\bar{\Psi}_{m}(s)$ represents a two-component spinor.
@@
II.2.2 Spin-orbit coupling term
In the narrow-tube limit an expression for the spin-orbit term was given in the Ref. Michaeli and Naaman (2019),
@@ II.2.2 Spin-orbit coupling term
As $s$ increases along the helix, $\mathbf{L}(s)$ precesses around the $z$-direction. This spatial dependency of ${\mathbf{L}}(s)$ is the main entry point of helicity into the quasi-one-dimensional model. The precession of ${\mathbf{L}}$ invites an analogy with a magnetic moment ${\bf\it\mu}$ in a rotating magnetic field with “Zeeman energy” $|{\bf\it\mu}\cdot\mathbf{B}|=\hbar\beta m$ and ${\bf B}=B\quantity[\sin(\Omega s)\hat{\mathbf{x}}-\cos(\Omega s)\hat{\mathbf{y}}+B_{z}/B\,\hat{\mathbf{z}}]$; $s$ being the effective “time” and oscillation frequency $\Omega=2\pi/\lambda\tilde{R}$. Motivated by this observation, in the following paragraph we adopt a transformation to the rotating frame Abragam (1961) in the next subsection.
II.2.3 Non-Abelian gauge transformation
@@
II.2.4 Minimal model Hamiltonian
@@
The $\tilde{\gamma}$-term corresponds to a momentum shift and thus represents a ”synthetic” vector potential. It can formally be removed from $\hat{H}^{\prime}_{N,m}$ by dressing the wavefunctions with a gauge-factor $e^{\mathrm{i}\tilde{\gamma}ms^{\prime}}$, Geyer, Gutierrez, and Cuniberti (2020) and therefore leaves the spectrum invariant. If $\tilde{\gamma}m$ has a natural interpretation as a vector potential, the term proportional to $b/R$ in the second line of (9) is the corresponding ”synthetic Zeeman term”. Notice that vector potential, $\tilde{\gamma}m$, and the Zeeman term change sign under $m\mapsto-m$; they do not break TRI because the full Hamiltonian sums over all $m$.
II.3 Dispersion relation
@@
The model (10) exhibits a two-band dispersion $E_{m,\alpha}(k)$, where $\alpha=\pm 1$; see supplementary material, Sec. S2, for explicit expressions.
@@
II.4 Spin and charge transport
II.4.1 Molecule bound to straight tubes
@@ II.4.2 Basic definitions
II.4.3 Transmissions and Landauer Formulæ
The Landauer formalism relates conductances to the spectral transmission probabilities: for the transmission from $\mathfrak{R}$ to $\mathfrak{L}$, $T^{(\mathfrak{L}\mathfrak{R})}_{\sigma\sigma^{\prime},m}(E)$, and vice versa, $T^{(\mathfrak{R}\mathfrak{L})}_{\sigma\sigma^{\prime},m}(E)$, and to the corresponding reflection amplitudes, $R^{(\mathfrak{L}\mathfrak{L})}_{\sigma\sigma^{\prime},m}(E)$ and $R^{(\mathfrak{R}\mathfrak{R})}_{\sigma\sigma^{\prime},m}(E)$; see Ref. Jacquod et al. (2012) for an overview. Here, the right (left) subscripts and superscripts of the transmission probability label the quantum numbers of an incoming (outgoing) wave, respectively. For example, $T^{(\mathfrak{R}\mathfrak{L})}_{\uparrow\downarrow,m}$ denotes the probability to transmit an electron with spin down from the left lead to the right one while flipping its spin in eigenchannel $m$. The charge conductances and the $z$-components of the spin conductances are given by
@@ II.4.3 Transmissions and Landauer Formulæ
where $E_{\mathrm{F}}$ indicates the Fermi energy (the supplementary material, Sec. S5, offers a standard derivation within the scattering formalism). The expression for $G$ chosen here emphasizes transmission of all spin species from left to right. From that expression the right spin conductance is obtained by multiplication by $\frac{1}{e}\frac{\hbar}{2}\sigma$ in the right lead. The left spin conductance expressed in the Eq. (12c) can then be understood as due to the reflected flux in the $\mathfrak{L}$-lead. Particle conservation (unitarity, see supplementary material, Sec. S4), leads to the equivalent form, Eq. (12d).
II.4.4 Transport Results
We address the transport problem by calculating the scattering matrix using conventional wavefunction matching. In this process, angular momentum $m$ matches at the two interfaces: it is conserved in the scattering process; for further details see supplementary material.
@@ II.4.4 Transport Results
  • Importantly, the spin conductances are non-zero for energies even far from the spin-orbit gap, where the oscillations peak at $\approx 0.2\,\frac{e}{4\pi}$. It can be shown that that $G_{x}^{(\mathfrak{L},\mathfrak{R})}(E_{\mathrm{F}})=0$ for any $E_{\mathrm{F}}$ since the Bloch functions have zero average $\hat{\sigma}_{x}$. Moreover, $G_{y}^{(\mathfrak{L},\mathfrak{R})}=0$ because the expectation values of $\hat{\sigma}_{y}$ change sign along with the sign change of $m$ (see Fig. 2), while the lead Hamiltonian is $m$-independent. In other words, the $y$ components of Bloch waves exactly cancel upon the summation over $m$.
  • Remarkably, the $\mathfrak{L}$ and $\mathfrak{R}$ spin conductances differ only up to a sign. By combining TRI and left-right reflection it is possible to derive relations between scattering matrix elements that lead to the exact identity $G_{z}^{(\mathfrak{R})}=-G_{z}^{(\mathfrak{L})}$ at any energy. (See supplementary material, Sec. S4,S5 for the details of this symmetry analysis.) In realistic molecular junctions a symmetry of couplings to the left and right leads can not be expected. We show in the Sec. II.6 that in this more general situation, the magnitude of $G_{z}^{(\mathfrak{R})}$ and $G_{z}^{(\mathfrak{L})}$ no longer is the same, while the sign is still opposite due to TRI.
Figure 3: Charge conductance and the $z$-components of the spin conductances in the Michaeli model as a function of the Fermi energy $E_{\mathrm{F}}$. Solid (dashed) lines are for a helix with 3 (6) turns, respectively . Parameters are the same as in Fig. 2, with a spin-orbit gap between 0.4 and 0.6.
II.5 Qualitative understanding of spin transport
II.5.1 A frequent misconception